# Listening to the Noise: Blind Denoising with Gibbs Diffusion

### v1

@@ 3.1 Denoising for Natural Images
We demonstrate the performance of our GDiff method on a noisy ImageNet test image (Fig. 3) characterized by $\sigma=0.2$ and $\varphi=0.4$. We run our Gibbs algorithm on $M=40$ iterations for 16 chains in parallel. We reach approximate convergence in less than 20 iterations. For our analysis, we consider only the final 20 samples from each chain. The figure presents, on the left, the original noisy image ${\bm{y}}$ alongside the true image ${\bm{x}}$, a denoised sample $\hat{{\bm{x}}}$, and the estimate of the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$ obtained by averaging all retained samples [^3] Note that the posterior mean estimate could have been alternatively computed using Tweedie’s formula (Robbins, 1956; Efron, 2011).. On the right, the inferred posterior distribution $q({\bm{\phi}}\,|\,{\bm{y}})$ is shown, demonstrating tight constraints around the true parameters and indicative of effective inference. Notably, the denoised posterior sample $\hat{{\bm{x}}}$ significantly enhances the image’s peak signal-to-noise ratio (PSNR). It is well known that optimal PSNR is attained for the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$ in the Bayesian picture. This is illustrated on this figure where the posterior mean estimate improves the PSNR of the sample by $\sim$3 dB. This is an interesting illustration of the antagonism between PSNR optimization and posterior sampling. The posterior sample is by construction a more realistic reconstruction in the light of the prior. Conversely, while the posterior mean yields a higher PSNR, indicating closer proximity to the true image, it may represent a less probable realization within the prior’s distribution. We provide in Fig. E.1 additional denoising examples.
@@
We validate our inference pipeline by focusing on the accuracy of the estimation of the noise parameters ${\bm{\phi}}=(\sigma,\varphi)$. We generate 800 noisy images using the ImageNet validation set, with noise parameters ${\bm{\phi}}$ sampled from their prior distribution $p({\bm{\phi}})$. For each of the noisy image ${\bm{y}}$, we conduct inferences using 4 parallel chains, each running for $M=60$ iterations. We discard the initial 30 samples as a warm-up phase. For each noisy observation, we compute the $\hat{R}$ convergence diagnostic and the effective sample size (ESS) (see e.g., Vehtari et al., 2021, and references therein). Fig. E.2 provides a scatter plot of these diagnostics across the parameter space. We find that the Markov chains achieve reasonable convergence, as indicated by $\hat{R}\lessapprox 1.1$, and that the ESS per chain is comparable to $M$.
@@ 3.2 Cosmological Inference from CMB Observations
The CMB closely approximates a Gaussian distribution, for which the covariance ${\bm{\Sigma}}_{\bm{\phi}}$ relates to cosmological models and ${\bm{\phi}}$ denotes the target cosmological parameters. Interestingly, component separation for CMB analysis can be viewed as a blind denoising problem of the form of Eq. (1) where the “noise” ${\bm{\varepsilon}}$ corresponds to the CMB, and the signal ${\bm{x}}$ represents the foregrounds. Following a rich history of dust foreground modeling efforts (e.g., Planck Collaboration XI, 2014; Allys et al., 2019; Vansyngel et al., 2017; Aylor et al., 2021; Regaldo-Saint Blancard et al., 2020, 2021; Thorne et al., 2021; Krachmalnicoff & Puglisi, 2021; Régaldo-Saint Blancard et al., 2023; Mudur & Finkbeiner, 2022) and going beyond the work of Heurtel-Depeiges et al. (2023), we focus here on a diffusion-based dust prior trained on simulations and apply GDiff for the Bayesian inference of cosmological parameters ${\bm{\phi}}$ given a mixture of CMB ${\bm{\epsilon}}$ and interstellar dust emission ${\bm{x}}$.
@@
We validate our inference pipeline similarly to Sect. 3.1. The $\hat{R}$ and ESS statistics (see Figs. E.3) for test observations generated across the entire parameter space, indicate that our sampler converges and mixes well. In addition, we conduct SBC for the parameters ${\bm{\phi}}$ and show in Fig. 7 the resulting rank statistics. These rank distributions are mostly compatible with the uniform distribution, indicating a well calibrated pipeline in particular for the inference of $H_{0}$ and $\omega_{b}$. However, for $\sigma$, the distribution is slanted, indicating bias for some of the chains. Here again, this must be the consequence of errors in the diffusion model, which impact the reconstructed power of the CMB estimates.
@@ Acknowledgements
[…]
@@ D.3 Existence of Stationary Distribution
More general definitions of compatibility can be found in references (Arnold et al., 2001; Liu et al., 2012). Sufficient and necessary conditions for compatibility are provided by Arnold et al. (2001, Theorem 1), restated here for convenience (and adapted to our notation).
@@
[…]
We find it enlightening to rewrite the last condition as an integral over the left-hand-side of Equation 25. Applying this condition to our problem we have
@@ D.4 Error in the Stationary Distribution
Appendix E Additional Results
@@
Figure E.1: Examples of blind denoising with GDiff on noisy ImageNet mocks with $\sigma=0.2$ and $\varphi\in\{-0.4,-0.2,0.0,0.2,0.4\}$ (from top to bottom). Left: Noisy examples ${\bm{y}}$ next to the noise-free images ${\bm{x}}$, denoised samples $\hat{{\bm{x}}}$ and estimates of the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$. Right: Inferred posterior distributions over the noise parameters.

### v2

@@ 3.1 Denoising for Natural Images
We demonstrate the performance of our GDiff method on a noisy ImageNet test image (Fig. 3) characterized by $\sigma=0.2$ and $\varphi=0.4$. We run our Gibbs algorithm on $M=40$ iterations for 16 chains in parallel. We reach approximate convergence in less than 20 iterations. For our analysis, we consider only the final 20 samples from each chain. The figure presents, on the left, the original noisy image ${\bm{y}}$ alongside the true image ${\bm{x}}$, a denoised sample $\hat{{\bm{x}}}$, and the estimate of the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$ obtained by averaging all retained samples [^3] Note that the posterior mean estimate could have been alternatively computed using Tweedie’s formula (Robbins, 1956; Efron, 2011).. On the right, the inferred posterior distribution $q({\bm{\phi}}\,|\,{\bm{y}})$ is shown, demonstrating tight constraints around the true parameters and indicative of effective inference. Notably, the denoised posterior sample $\hat{{\bm{x}}}$ significantly enhances the image’s peak signal-to-noise ratio (PSNR). It is well known that optimal PSNR is attained for the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$ in the Bayesian picture. This is illustrated on this figure where the posterior mean estimate improves the PSNR of the sample by $\sim$3 dB. This is an interesting illustration of the antagonism between PSNR optimization and posterior sampling. The posterior sample is by construction a more realistic reconstruction in the light of the prior. Conversely, while the posterior mean yields a higher PSNR, indicating closer proximity to the true image, it may represent a less probable realization within the prior’s distribution. We provide in Fig. F.1 additional denoising examples.
@@
We validate our inference pipeline by focusing on the accuracy of the estimation of the noise parameters ${\bm{\phi}}=(\sigma,\varphi)$. We generate 800 noisy images using the ImageNet validation set, with noise parameters ${\bm{\phi}}$ sampled from their prior distribution $p({\bm{\phi}})$. For each of the noisy image ${\bm{y}}$, we conduct inferences using 4 parallel chains, each running for $M=60$ iterations. We discard the initial 30 samples as a warm-up phase. For each noisy observation, we compute the $\hat{R}$ convergence diagnostic and the effective sample size (ESS) (see e.g., Vehtari et al., 2021, and references therein). Fig. F.2 provides a scatter plot of these diagnostics across the parameter space. We find that the Markov chains achieve reasonable convergence, as indicated by $\hat{R}\lessapprox 1.1$, and that the ESS per chain is comparable to $M$.
@@ 3.2 Cosmological Inference from CMB Observations
The CMB closely approximates a Gaussian distribution, for which the covariance ${\bm{\Sigma}}_{\bm{\phi}}$ relates to cosmological models and ${\bm{\phi}}$ denotes the target cosmological parameters. Interestingly, component separation for CMB analysis can be viewed as a blind denoising problem of the form of Eq. (1) where the “noise” ${\bm{\varepsilon}}$ corresponds to the CMB, and the signal ${\bm{x}}$ represents the foregrounds. Following a rich history of dust foreground modeling efforts (e.g., Planck Collaboration XI, 2014; Allys et al., 2019; Vansyngel et al., 2017; Aylor et al., 2021; Regaldo-Saint Blancard et al., 2020, 2021; Thorne et al., 2021; Krachmalnicoff & Puglisi, 2021; Régaldo-Saint Blancard et al., 2023; Mudur & Finkbeiner, 2022) and going beyond the work of Heurtel-Depeiges et al. (2023), we focus here on a diffusion-based dust prior trained on simulations and apply GDiff for the Bayesian inference of cosmological parameters ${\bm{\phi}}$ given a mixture of CMB ${\bm{\varepsilon}}$ and interstellar dust emission ${\bm{x}}$.
@@
We validate our inference pipeline similarly to Sect. 3.1. The $\hat{R}$ and ESS statistics (see Figs. F.3) for test observations generated across the entire parameter space, indicate that our sampler converges and mixes well. In addition, we conduct SBC for the parameters ${\bm{\phi}}$ and show in Fig. 7 the resulting rank statistics. These rank distributions are mostly compatible with the uniform distribution, indicating a well calibrated pipeline in particular for the inference of $H_{0}$ and $\omega_{b}$. However, for $\sigma$, the distribution is slanted, indicating bias for some of the chains. Here again, this must be the consequence of errors in the diffusion model, which impact the reconstructed power of the CMB estimates.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the fields of Denoising and Machine Learning for Science. There are potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
${\bm{z}}_{t}=a(t){\bm{x}}+b(t){\bm{\varepsilon^{\prime}}},\hfill\text{with}~{}\varepsilon^{\prime}\sim\mathcal{N}(0,{\bm{\Sigma}}_{\bm{\varphi}}),$ (11)
@@ D.3 Existence of Stationary Distribution
More general definitions of compatibility can be found in references (Arnold et al., 2001; Liu et al., 2012). Sufficient conditions for compatibility are provided by Arnold et al. (2001, Theorem 1), restated here for convenience (and adapted to our notation). [^9] As argued by Liu et al. (2021, Theorem 2.3), the conditions by Arnold et al. (2001) are not necessary and item 1 in Theorem D.2 can be replaced with a weaker condition.
@@
    $\int_{\mathcal{X}}u({\bm{x}})\text{d}{\bm{x}}<\infty.$ (25)
We find it enlightening to rewrite the last condition as an integral over the left-hand-side of Eq. (25). Applying this condition to our problem we have
@@ D.4 Error in the Stationary Distribution
Appendix E Limitations of Guidance-based Posterior Sampling Methods
By examining the applicability of existing methods DPS (Chung et al., 2023) and $\Pi$-GDM (Song et al., 2023a) in the cosmology application presented in Sect. 3.2, we identified limitations in these guidance-based methods, which further legitimize the interest in our approach. We report our investigation in this appendix.
First, it is important to note that, since the exact conditional guidance term is typically intractable, guidance-based posterior sampling methods require methodological approximations. These approximations may impact the estimation of the posterior distribution in unpredictable ways. For scientific applications like our problem of cosmological inference from CMB data, this can be a significant downside. Furthermore, we encountered significant numerical difficulties with DPS and $\Pi$-GDM, which we consider strong evidence of the limitations of these methods for our cosmology application.
Let us first rewrite Eq. (1) as a standard linear inverse problem equation (i.e., involving white Gaussian noise). Starting from Eq. (1):
${\bm{y}}={\bm{x}}+{\bm{\varepsilon}},\quad\text{with}\;{\bm{\varepsilon}}\sim\mathcal{N}({\bm{0}},{\bm{\Sigma}}_{\bm{\phi}}),$ (31)
and assuming ${\bm{\Sigma}}_{\bm{\phi}}$ is invertible, we define $\tilde{{\bm{\varepsilon}}}={\bm{\Sigma}}_{\bm{\phi}}^{-1/2}{\bm{\varepsilon}}$, so that $\tilde{{\bm{\varepsilon}}}\sim\mathcal{N}({\bm{0}},{\bm{I}}_{d})$. Then, introducing $\tilde{{\bm{y}}}_{{\bm{\phi}}}={\bm{\Sigma}}_{\bm{\phi}}^{-1/2}{\bm{y}}$ and ${\bm{A}}_{{\bm{\phi}}}={\bm{\Sigma}}_{\bm{\phi}}^{-1/2}$, we get:
$\tilde{{\bm{y}}}_{{\bm{\phi}}}={\bm{A}}_{{\bm{\phi}}}{\bm{x}}+\tilde{{\bm{\varepsilon}}},$ (32)
which resembles a standard linear inverse problem where the linear operator is ${\bm{A}}_{{\bm{\phi}}}={\bm{\Sigma}}_{\bm{\phi}}^{-1/2}$. Also, note that the observation $\tilde{{\bm{y}}}_{{\bm{\phi}}}$ depends on the parameters ${\bm{\phi}}$, which adds a layer of complexity when the goal is to infer ${\bm{\phi}}$.
We implemented off-the-shelf DPS (Chung et al., 2023) and $\Pi$-GDM (Song et al., 2023a) algorithms in the context of our cosmology application [^10] Note that, in our case, the $\Pi$-GDM guidance term is straightforward to compute since the operator ${\bm{\Sigma}}_{\bm{\phi}}^{-1/2}$ is diagonal in Fourier space.. The standard regularization (guidance strength) schemes initially yielded numerical float overflows. To avoid these numerical limitations, we had to decrease the guidance strength to a point where it was clear that we were not sampling the posterior distribution (large scales structure were off, as shown in Fig. E.1). Our most satisfying approach turned out to be a variation of $\Pi$-GDM involving a time-dependent regularization term (Song et al., 2023a, App. A.3) way outside the normal range for any inverse problems. Fig. E.2 depicts our results with associated summary statistics, proving that the posterior distribution is very poorly estimated. Furthermore, when trying to use this last method in our Gibbs pipeline, we could not reach convergence to any posterior distribution over $\phi$.
Figure E.1: Best results when optimizing the guidance strength for the standard off-the shelf $\Pi$-GDM algorithm.
Figure E.2: Best results when optimizing a regularization in a modified $\Pi$-GDM.
We interpret these numerical challenges as the result of the significant dynamic range of eigenvalues of ${\bm{\Sigma}}_{\bm{\phi}}$. For our cosmology application, the condition number of ${\bm{\Sigma}}_{\bm{\phi}}$ is approximately $6\times 10^{5}$, and the largest eigenvalue of ${\bm{\Sigma}}_{\bm{\phi}}^{-1/2}$ is $1.6\times 10^{2}$. In comparison, standard inverse problems such as inpainting, deblurring, or super-resolution problems, typically involve operators with eigenvalues bounded between 0 and 1. Since guidance-based methods rely on estimates of $\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}_{t}|{\bm{y}},{\bm{\phi}})$, these large eigenvalues lead to numerical instabilities along their corresponding eigenvectors (i.e., for our problem, at high frequencies). If we attempt to mitigate this instability using regularization, the guidance along eigenvectors linked to the smallest eigenvalues of ${\bm{\Sigma}}_{\bm{\phi}}^{-1/2}$ becomes insufficient, resulting in impaired low frequencies/large-scale structures.
Appendix F Additional Results
@@
Figure F.1: Examples of blind denoising with GDiff on noisy ImageNet mocks with $\sigma=0.2$ and $\varphi\in\{-0.4,-0.2,0.0,0.2,0.4\}$ (from top to bottom). Left: Noisy examples ${\bm{y}}$ next to the noise-free images ${\bm{x}}$, denoised samples $\hat{{\bm{x}}}$ and estimates of the posterior mean $\mathbb{E}\left[{\bm{x}}\,|\,{\bm{y}}\right]$. Right: Inferred posterior distributions over the noise parameters.